# Towards Single-Lens Controllable Depth-of-Field Imaging via All-in-Focus Aberration Correction and Monocular Depth Estimation

### v1

@@ I Introduction
In photography, an All-in-Focus (AiF) image may not effectively convey the creator’s intent. Professional photographers often use Depth of Field (DoF), which refers to the range of depths in a scene that is imaged sharply in focus, as a composition tool to achieve amazing artistic effects. This range is mainly determined by the aperture of the camera lens used for shooting: a narrow aperture produces a wide DoF, and a wide aperture produces a shallow DoF [1]. In portraiture, for example, a shallow DoF and strong background bokeh enable the photographer to highlight a subject. However, achieving controllable DoF requires a well-designed high-end lens, which is costly, inconvenient, and often challenging to operate, limiting their widespread use in low-budget scenarios or on low-payload platforms [2, 3]. An interesting question arises: Is it feasible to achieve controllable DoF imaging using only a Minimalist Optical System (MOS), such as a single-lens optical system? Due to the lack of necessary optical optimization variables, MOS suffers from low imaging quality and an inflexible DoF. Yet, the scene information can be reconstructed by predicting the AiF aberration-free image and the corresponding depth map of the aberration image, making our goal attainable. Recently, the great success of deep learning-based methods have promoted the progress of Monocular Depth Estimation (MDE) [4, 5, 6, 7, 8, 9, 10, 11] and Computational Aberration Correction (CAC) [12, 13, 14, 15, 16, 17].
@@
For obtaining scene information, UniDepth [11], as a generalized MDE model, is used to generate reliable depth maps. To aid in recovering the aberration-free AiF image from the depth-aware optical degradation, we establish the Depth-aware Aberration MOS (DAMOS) dataset by ray tracing the PSF map of different object distances and introduce the Depth-aware Degradation-adaptive Training (DA2T) scheme. Considering that depth information may enable the network to understand the optical degradation process that varies with depth, we design two plug-and-play depth-aware mechanisms: Residual Depth-Image Cross-Attention Block (RDICAB) and Depth-aware Deformable Convolution Block (D2CB). In RDICAB, the cross-attention mechanism is utilized for the initial fusion of depth features and image features, while in D2CB, the image feature is modulated by the depth feature through deformable convolution to adaptively learn the depth-aware degradation.
@@
To assess the effectiveness of our proposed DA2T scheme and two depth-aware mechanisms, we conduct extensive experiments to test the generalizability of the trained CAC model on various datasets, including both simulated and real-world datasets. The experimental results demonstrate that the model trained by the DA2T scheme generalizes better than the non-DA2T restoration model in terms of CAC. Specifically, the DA2T model successfully restores scenes at different depths, whereas the non-depth-aware model fails to do so. Results also prove that under the same DA2T scheme, both RDICAB and D2CB enable the model to restore more details and reduce artifacts by fusing depth information. To further verify the generalizability of our method, we demonstrate results on four state-of-the-art Super-Resolution (SR) architectures [23, 24, 25, 26], on two single-lens systems (MOS-S1 and MOS-S2), showing that our approach is not strictly tied to a specific restoration module and can consistently enhance the recovery performance.
The experimental results show that high-quality AiF aberration correction is achieved, providing a favorable precondition for controllable DoF imaging. By utilizing the restored image, the estimated depth map, and the PSF map inferred from the Omni-Lens-Field, controllable DoF imaging with multiple lenses is achieved.
@@ II-B Computational Aberration Correction
Most classic restoration methods [40, 41, 42, 43, 44] attribute image degradation to a convolution process. In this way, they restore degraded data by estimating the degraded kernel. However, the effectiveness of these model-based methods relies on the accuracy of the estimated degraded kernel and suffers from poor generalization performance. In the last decade, deep-learning-based methods have been widely employed in image restoration [45, 46, 47], and the CAC area has also made great progress, achieving remarkable restoration results. These methods typically use an encoder-decoder network structure to fit the degradation process of the optical system through reverse gradient optimization [22, 14]. Diverging from purely data-driven algorithms, some studies [48, 15, 16, 49, 28] investigate the integration of spatial-variant optical prior knowledge, such as the FoV and PSF cues, into the network to enhance image restoration.
Different from these works, to further enhance the recovery performance, we pioneer using scene-depth information to aid in CAC for addressing depth-varying optical degradation. We design two plug-and-play depth-aware mechanisms (RDICAB and D2CB) that can effectively improve recovery performance.
@@ II-C Controllable Depth-of-Field Imaging
Earlier work on DoF manipulation is achieved through hardware [50, 51, 52, 53]. Hong et al. [52] propose a method using a spatial light modulator at the pupil plane to control the DoF in optical microscopes flexibly. Kuthirumma et al. [53] design a camera system with flexible DoF by controlling the motion of the detector during the image integration process. The rapid advancements in deep learning have led to the emergence of a growing number of learning-based approaches for controlling DoF. With the availability of an input or an estimated depth map, a shallow DoF image can be synthesized through either a neural network-based approach [54, 55] or a hybrid method combining classical and neural rendering techniques [56]. In addition, one method to achieve controllable DoF imaging is to capture a focus stack, and then composite the relevant frames to emulate the target DoF, but this approach suffers from the lengthy capture time, which restricts its use to static scenes [57]. Alzayer et al [58] leverage the dual-camera system on the smartphone to achieve various DoF manipulations through defocus control. However, the optical systems considered in the aforementioned methods are idealized, without accounting for the potential image degradation caused by optical aberrations. Therefore, in our work, the proposed DCDI framework enables controllable DoF imaging for the MOS suffering from aberration-induced degradation and uncontrolled DoF. It is essential to emphasize that the proposed method can synthesize the imaging effects of any lens with known parameters and control its DoF.
@@ III-A Depth-aware Image Dataset Construction
In this part, we introduce how to build a depth-aware aberration image dataset, including the construction of the PSFLib and image simulation.
@@
Here, $r_{m,n}^{i,j}$ represents the distance between the ray with pixel index $(m,n)$ and the central ray of the image plane, where $(i,j)$ is the index of the ray in the pupil. We trace $128{\times}128$ rectangular distributed rays in the pupil in practice. $\sigma{=}\sqrt{\Delta h^{2}{+}\Delta w^{2}}/3$, and ${\Delta}h{\times}{\Delta}w$ is the physical size of each pixel.
After obtaining the PSFs for the sampled field of view and wavelength, we can then synthesize the three-channel RGB PSFs by simulating the spectral response curve of the CMOS sensor [59]. The three-channel RGB PSF for the same imaging depth can be obtained by the following formula:
@@
where $c$ represents one of R, G, and B channels, $\theta$ represents the sampled field of view, which is obtained by dividing the maximum image height of the optical system into $N_{fov}$ parts, $d$ represents the imaging depth, and $R(\lambda_{s})$ represents normalized wavelength response coefficient. To trade off the accuracy and speed of the simulation, the $N_{fov}$ is set to $64$, and the RGB PSF size is set to $3{\times}41{\times}41$. The common optical systems used in industry, including our MOS, are axisymmetric, resulting in PSFs that don’t vary with azimuth. Therefore, the PSF map can be efficiently filled by interpolating, rotating, and resizing $PSF(c,\theta,d)$. The RGB PSF map at a scene depth can be described as:
@@
where $\ast$ represents the convolution operation, $I_{H}$ and $I_{L}$ denote ground truth and aberration image respectively, $\psi$ denotes the searching function that queries the PSFLib and outputs the corresponding PSF of every patch with different scene depths. The depth value of each patch is the average depth value for all pixels within that patch. The average depth map of the entire scene can be expressed as follows
@@ III-B Depth-aware Aberration Correction Network
As demonstrated in [16], the SR paradigm achieves better results in aberration image recovery. For this reason, the proposed Depth-aware Aberration Correction (DACN) is an end-to-end CAC model based on the SR paradigm, enabling deep multimodal fusion of image and depth information. The architecture of the proposed model is shown in Fig. 3, which includes our proposed two plug-and-play depth-aware mechanisms. Residual Depth-Image Cross-Attention Block (RDICAB) follows the shallow feature extraction layer. It is employed for the initial stage of cross-modal fusion, aiming to integrate image features and deep features. Depth-aware Deformable Convolution Block (D2CB) is employed within each Depth-aware Aberration Correction Block (DACB). The image features are modulated by the depth features through a Deformable Convolution Layer (DCL) [61], thereby effectively learning the depth-aware aberration degradation process.
@@
where $(\cdot)^{T}$ denotes the transpose operator. The output of the cross-attention operation is added to $I_{i}$, and then the result of the addition is processed by a Multi-Layer Perceptron (MLP) and a $1{\times}1$ convolution layer respectively. Finally, a Long Skip Residual Connection (LSRC) is used to add $I_{i}$ to the intermediate features to obtain final output fusion features $F_{o}$. Overall, RDICAB, utilizing the cross-attention mechanism, effectively handles the initial fusion of multimodal features and provides high-quality modulated image features for subsequent modules.
Depth-aware Deformable Convolution Block (D2CB). Different from standard Deformable Convolutional Networks (DCN) [64], which calculates offsets based on image feature maps of the same modality, we only calculate offsets through depth features to modulate image features, thereby guiding the network to learn the depth-variant degradation process. As shown in the lower right of Fig. 3, D2CB consists of a DCL [61] and an offset net. The modulated output image feature $I_{o}$ is generated by a DCL that takes the input image feature $I_{i}$ and an offset produced by feeding the depth feature $D_{i}$ into an offset network. Modulated output $I_{o}$ is formulated as:
@@ IV-A Datasets
Synthetic datasets. We establish a Depth-aware Aberration MOS (DAMOS) dataset to train our CAC model, supporting AiF aberration correction. The absence of datasets including depth-aware aberration images, corresponding depth maps, and AiF images, prompts us to focus on datasets for CAC consisting of AiF images and depth ground truth and to generate depth-aware aberration images by blurring the AiF images. Therefore, we simulate the 4D PSFs on two MOS with distinct aberration distribution (MOS-S1 and MOS-S2) and then use it to perform patch-wise convolution on each color channel of the high-quality AiF images and reunite them to obtain the corresponding synthetic aberration image. We use the NYUv2 dataset [67], which comprises $1,449$ images of the resolution of $480{\times}640$ as the high-quality AiF images to generate training and test data pairs. The Middlebury2014 dataset [68] is also used as the testing dataset, which includes $23$ high-resolution real-world RGB-D images. The sensor resolution determines the resolution ratio of training and testing images, and the aberration property of the lens can be altered by any resizing operation. However, the extremely high-resolution images of $4000{\times}6000$ pixels produced by the Sony ${\alpha}$6600 camera sensor cannot be handled during network training processes. Due to the physical properties of the PSF, downsampling is feasible [66, 28]. Therefore, we downsample the sensor to a $1280{\times}1920$ pixel size considering the resolution of the used datasets. For the NYUv2 dataset, we treat the images as part of a $1280{\times}1920$ resolution image to enable the model trained on this dataset to restore the real-world data without any finetuning. The operation process is illustrated in the Appendix. The resolution of the images in Middlebury2014 is reshaped and cropped to $1280{\times}1920$. Following the setup in [15], we randomly perturb the parameters of the MOS within a range of $10\%$ to generate image pairs with different aberration distributions, which can alleviate the domain gap problem between the synthetic and real-world images.
@@
Figure 6: Visualization of PSFs across varying FoVs (normalized diagonal field $fov{\in}\left[0,1\right]$) and scene depths (depth $d{\in}\left[0.7m,10m\right]$ and depth $d{\geq}10m$ is regarded as infinity). The PSFs of MOS-S1 and MOS-S2 are both spatially non-uniform and exhibit depth-dependent variations. The MOS-S1 has a focal length of $20.0mm$, an F-number of $5.0$, and a FoV of $42.6^{\circ}$. The MOS-S2 has a focal length of $24.7mm$, an F-number of $3.5$, and a FoV of $34.0^{\circ}$.

### v2

# Towards Single-Lens Controllable Depth-of-Field Imaging via Depth-Aware Point Spread Functions

@@ I Introduction
In photography, an All-in-Focus (AiF) image may not effectively convey the creator’s intent. Professional photographers often use Depth of Field (DoF), which refers to the range of depths in a scene that is imaged sharply in focus, as a composition tool to achieve amazing artistic effects. This range is mainly determined by the aperture of the camera lens used for shooting: a narrow aperture produces a wide DoF, and a wide aperture produces a shallow DoF [1]. In portraiture, for example, a shallow DoF and strong background bokeh enable the photographer to highlight a subject. However, achieving controllable DoF requires a well-designed high-end lens, which is costly, inconvenient, and often challenging to operate, limiting their widespread use in low-budget scenarios or on low-payload platforms [2, 3]. An interesting question arises: Is it feasible to achieve controllable DoF imaging using only a Minimalist Optical System (MOS), such as a single-lens optical system? Due to the lack of necessary optical optimization variables, MOS suffers from low imaging quality and an inflexible DoF. Yet, the scene information can be reconstructed by predicting the AiF aberration-free image and the corresponding depth map of the aberration image, making our goal attainable. Recently, the great success of deep-learning-based methods have promoted the progress of Monocular Depth Estimation (MDE) [4, 5, 6, 7, 8, 9, 10, 11] and Computational Aberration Correction (CAC) [12, 13, 14, 15, 16, 17].
@@
For obtaining scene information, a generalized metric MDE model [11] is used to generate reliable metric depth maps. To aid in recovering the aberration-free AiF image from the depth-aware optical degradation, we establish the Depth-aware Aberration MOS (DAMOS) dataset by ray tracing the PSF map of different object distances and introduce the Depth-aware Degradation-adaptive Training (DA2T) scheme. Considering that depth information may enable the network to understand the optical degradation process that varies with depth, we design two plug-and-play depth-aware mechanisms: Residual Depth-Image Cross-Attention Block (RDICAB) and Depth-aware Deformable Convolution Block (D2CB). In RDICAB, the cross-attention mechanism is utilized for the initial fusion of depth features and image features, while in D2CB, the image feature is modulated by the depth feature through deformable convolution to adaptively learn the depth-aware degradation.
@@
To assess the effectiveness of our proposed DA2T scheme and two depth-aware mechanisms, we conduct extensive experiments to test the generalizability of the trained CAC model on various datasets, including both simulated and real-world datasets. The quantitative and qualitative experimental results demonstrate that the model trained by the DA2T scheme generalizes better than the non-DA2T restoration model in terms of CAC. Specifically, the DA2T model successfully restores scenes at different depths, whereas the non-depth-aware model fails to do so. The quantitative and qualitative results also prove that under the same DA2T scheme, both RDICAB and D2CB enable the model to restore more details and reduce artifacts by fusing depth information. To further verify the generalizability of our method, we demonstrate results on four state-of-the-art Super-Resolution (SR) architectures [23, 24, 25, 26], on two single-lens systems (MOS-S1 and MOS-S2), showing that our approach is not strictly tied to a specific restoration module and can consistently enhance the recovery performance.
The experimental results show that high-quality AiF aberration correction is achieved, providing a favorable precondition for controllable DoF imaging. By utilizing the restored image, the estimated depth map, and the PSF map inferred from the Omni-Lens-Field, controllable DoF imaging with multiple lenses is achieved. Considering the lack of Ground Truth (GT) for the controllable DoF imaging, we only present “qualitative” results instead of conducting the GT-required “quantitative” evaluation.
To enhance readability, Table I presents all abbreviations and their full forms used in this paper.
TABLE I: Glossary of Terms and Abbreviations
Abbreviation Full Form
AiF All in Focus
CAC Computational Aberration Correction
CoC Circle-of-Confusion
DAMOS Depth-aware Aberration MOS
DA2T Depth-aware Degradation-adaptive Training
DACN Depth-aware Aberration Correction Network
D2CB Depth-aware Deformable Convolution Block
DCDI Depth-aware Controllable DoF Imaging
DCL Deformable Convolution Layer
DoF Depth-of-Field
FoV Field of View
K Key
MDE Monocular Depth Estimation
MLP Multi-Layer Perceptron
MOS Minimalist Optical System
PSF Point Spread Function
PSFLib PSF Library
PSA Permuted Self-Attention
Q Query
RB Residual Block
RDG Residual Deep-feature-extraction Group
RDICAB Residual Depth-Image Cross-Attention Block
RSTB Residual Swin Transformer Block
V Value
@@ II-B Computational Aberration Correction
Most classic restoration methods [40, 41, 42, 43, 44] attribute image degradation to a convolution process. In this way, they restore degraded data by estimating the degraded kernel. However, the effectiveness of these model-based methods relies on the accuracy of the estimated degraded kernel and suffers from poor generalization performance. In the last decade, deep-learning-based methods have been widely employed in image restoration [45, 46, 47], and the CAC area has also made great progress, achieving remarkable restoration results. These methods typically use an encoder-decoder network structure to fit the degradation process of the optical system through reverse gradient optimization [22, 14]. Diverging from purely data-driven algorithms, some studies [48, 15, 16, 49, 28] investigate the integration of spatial-variant optical prior knowledge, such as the FoV and PSF cues, into the network to enhance image restoration. Additionally, some researches [50, 51] have proposed defocus deblurring methods that account for depth-varying PSFs; however, they often overlook the aberrations introduced by imperfections in the optical system design.
To address the degradation caused by real-world lenses, Luo et al. [28] estimated the depth map of the aberration image and selected the corresponding PSF from PSFLib to restore the aberration image. However, their method did not directly use depth information for aberration correction or other tasks. In contrast, our framework leverages the estimated depth from the aberration image to enhance CAC performance, addressing depth-aware degradation and enabling effective restoration of scenes with varying depths. Moreover, we explore the use of estimated depth maps and restored images for other applications, such as controllable DoF imaging.
@@ II-C Controllable Depth-of-Field Imaging
Earlier work on DoF manipulation is achieved through hardware [52, 53, 54, 55]. Hong et al. [54] propose a method using a spatial light modulator at the pupil plane to control the DoF in optical microscopes flexibly. Kuthirumma et al. [55] design a camera system with flexible DoF by controlling the motion of the detector during the image integration process. The rapid advancements in deep learning have led to the emergence of a growing number of learning-based approaches for controlling DoF. With the availability of an input or an estimated depth map, a shallow DoF image can be synthesized through either a neural network-based approach [56, 57] or a hybrid method combining classical and neural rendering techniques [58]. In addition, one method to achieve controllable DoF imaging is to capture a focal stack, and then composite the relevant frames to emulate the target DoF, but this approach suffers from the lengthy capture time, which restricts its use to static scenes [59]. Alzayer et al. [60] leverage the dual-camera system on the smartphone to achieve various DoF manipulations through defocus control. However, the optical systems considered in the aforementioned methods are idealized, without accounting for the potential image degradation caused by optical aberrations. Therefore, in our work, the proposed DCDI framework enables controllable DoF imaging for the MOS suffering from aberration-induced degradation and uncontrolled DoF. It is essential to emphasize that the proposed method can synthesize the imaging effects of any lens with known parameters and control its DoF.
@@ III-A Depth-aware Image Dataset Construction
In this part, we introduce how to build a depth-aware aberration image dataset, including the construction of the PSFLib and image simulation. The construction of the depth-aware image dataset, as a key component, is integrated into the training process with the aim of embedding depth-varying optical characteristics into network training. A detailed introduction to DA2T is provided in the appendix.
@@
Here, $r_{m,n}^{i,j}$ represents the distance between the ray with pixel index $(m,n)$ and the central ray of the image plane, where $(i,j)$ is the index of the ray in the pupil.
We trace $128{\times}128$ rectangular distributed rays in the pupil in practice. $\sigma{=}\sqrt{\Delta h^{2}{+}\Delta w^{2}}{/}3$, and ${\Delta}h{\times}{\Delta}w$ is the physical size of each pixel. In the standard deviation formula, the number $3$ is chosen as the magic number based on the $3{\sigma}$ rule, which ensures that $99.7\%$ of the light energy is concentrated within ±${\pm}\sigma$, allowing us to limit the calculation area to one pixel rather than integrating over an infinite range. This method greatly reduces the computational complexity while maintaining the high accuracy of the results. After obtaining the PSFs for the sampled field of view and wavelength, we can then synthesize the three-channel RGB PSFs by simulating the spectral response curve of the CMOS sensor [61]. The three-channel RGB PSF for the same imaging depth can be obtained by the following formula:
@@
where $c$ represents one of R, G, and B channels, $\theta$ represents the sampled field of view, which is obtained by dividing the maximum image height of the optical system into $N_{fov}$ parts, $d$ represents the imaging depth, and $R(\lambda_{s})$ represents normalized wavelength response coefficient. It is worth noting that while the focal length does affect the PSF as it represents the optical system’s ability to converge light, it is a fixed parameter for a given fixed-focus optical system. In Eq. 3, we aim to emphasize the impact of variations in Field of View (FoV), channel, and object distance on the PSF under the condition of a fixed focal length. Therefore, it is not explicitly expressed in the equation. To trade off the accuracy and speed of the simulation, the $N_{fov}$ is set to $64$, and the RGB PSF size is set to $3{\times}41{\times}41$. The common optical systems used in industry, including our MOS, are axisymmetric, resulting in PSFs that don’t vary with azimuth. Therefore, the PSF map can be efficiently filled by interpolating, rotating, and resizing $PSF(c,\theta,d)$. The RGB PSF map at a scene depth can be described as:
@@
where $\ast$ represents the convolution operation, $I_{H}$ and $I_{L}$ denote ground truth and aberration image respectively, $\psi$ denotes the searching function that queries the PSFLib and outputs the corresponding PSF of every patch with different scene depths. Here, we also consider the distortion of the optical lens, and perform distortion processing on both the aberration image and the clear image to ensure pixel alignment. Since distortion does not affect the image quality, distortion correction is not within the scope of this paper. The depth value of each patch is the average depth value for all pixels within that patch. The average depth map of the entire scene can be expressed as follows
@@ III-B Depth-aware Aberration Correction Network
As demonstrated in [16], the SR paradigm achieves better results in aberration image recovery. It should be emphasized that the image resolution is not improved; instead, the pixel unshuffle operation is applied to the input images and depth maps to ensure consistency between input and output resolutions. For this reason, the proposed Depth-aware Aberration Correction (DACN) is an end-to-end CAC model based on the SR paradigm, enabling deep multimodal fusion of image and depth information. The architecture of the proposed model is shown in Fig. 3, which includes our proposed two plug-and-play depth-aware mechanisms. Residual Depth-Image Cross-Attention Block (RDICAB) follows the shallow feature extraction layer. It is employed for the initial stage of cross-modal fusion, aiming to integrate image features and deep features. Depth-aware Deformable Convolution Block (D2CB) is employed within each depth-aware aberration correction block. The image features are modulated by the depth features through a Deformable Convolution Layer (DCL) [63], which effectively facilitates the learning of the depth-aware aberration degradation process.
@@
where $(\cdot)^{T}$ denotes the transpose operator. The output of the cross-attention operation is added to $I_{i}$, and then the result of the addition is processed by a Multi-Layer Perceptron (MLP) and a $1{\times}1$ convolution layer. Finally, a long skip residual connection is used to add $I_{i}$ to the intermediate features to obtain final output fusion features $F_{o}$. Overall, RDICAB, utilizing the cross-attention mechanism, effectively handles the initial fusion of multimodal features and provides high-quality modulated image features for subsequent modules.
Depth-aware Deformable Convolution Block (D2CB). Due to the irregular shape and size of spatial-variant PSFs, traditional convolution operations, which rely on fixed positional features around the center, are unable to effectively handle these degradations. In contrast, deformable convolution can adapt the shape of the convolution kernel by introducing offsets. This flexibility allows depth information, which encompasses the characteristics of spatial variation, to guide the convolution process, enhancing the model’s ability to handle depth-aware optical degradation. Different from standard deformable convolutional networks [66], which calculates offsets based on image feature maps of the same modality, we only calculate offsets through depth features to modulate image features. As shown in the lower right of Fig. 3, D2CB consists of a DCL [63] and an offset network. The modulated output image feature $I_{o}$ is generated by a DCL that takes the input image feature $I_{i}$ and an offset produced by feeding the depth feature $D_{i}$ into an offset network. The modulated output $I_{o}$ is formulated as:
@@ IV-A Datasets
Synthetic datasets. We establish a Depth-aware Aberration MOS (DAMOS) dataset to train our CAC model, supporting AiF aberration correction. The absence of datasets including depth-aware aberration images, corresponding depth maps, and AiF images, prompts us to focus on datasets for CAC consisting of AiF images and depth ground truth and to generate depth-aware aberration images by blurring the AiF images. Therefore, we simulate the 4D PSFs on two MOS with distinct aberration distribution (MOS-S1 and MOS-S2) and then use it to perform patch-wise convolution on each color channel of the high-quality AiF images and reunite them to obtain the corresponding synthetic aberration image. We use the NYUv2 dataset [69], which comprises $1,449$ images of the resolution of $480{\times}640$ as the high-quality AiF images to generate training and test data pairs. The Middlebury2014 dataset [70] is also used as the testing dataset, which includes $23$ high-resolution real-world RGB-D images. The sensor resolution determines the resolution ratio of training and testing images, and the aberration property of the lens can be altered by any resizing operation. However, the extremely high-resolution images of $4000{\times}6000$ pixels produced by the Sony ${\alpha}$6600 camera sensor cannot be handled during network training processes. Due to the physical properties of the PSF, downsampling is feasible [68, 28]. Therefore, we downsample the sensor to a $1280{\times}1920$ pixel size using area interpolation, which preserves more image details and reduces the generation of artifacts, considering the resolution of the used datasets. For the NYUv2 dataset, we treat the images as part of a $1280{\times}1920$ resolution image to enable the model trained on this dataset to restore the real-world data without any finetuning. The resolution of the images in Middlebury2014 is reshaped and cropped to $1280{\times}1920$. Following the setup in [15], we randomly perturb the parameters of the MOS within a range of $10\%$ to generate image pairs with different aberration distributions, which can alleviate the domain gap between synthetic and real-world images.
@@
Figure 6: Visualization of PSFs across varying FoVs (normalized diagonal field $fov~{}{\in}~{}\left[0,1\right]$) and scene depths (depth $d~{}{\in}~{}\left[0.7~{}m,10~{}m\right]$ and depth $d~{}{\geq}~{}10m$ is regarded as infinity). The PSFs of MOS-S1 and MOS-S2 are both spatially non-uniform and exhibit depth-dependent variations. The MOS-S1 has a focal length of $20.0~{}mm$, an F-number of $5.0$, and a FoV of $42.6^{\circ}$. The MOS-S2 has a focal length of $24.7~{}mm$, an F-number of $3.5$, and a FoV of $34.0^{\circ}$. The size of each PSF is $41{\times}41$ at a pixel size of $12.2~{}{\mu}m$.